• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reitz

    Log entry time 16:05:09 on March31,2005

    Entry number 142890

    keyword=Day Shift Summary

    8:00 took over from shift, they reported having problems
    with ROC 5 and not seeing ADC histograms for BigBite.
    Looking at ROC 5, we see error messages from QDC ID2 and ID3.
    Reboot does not help, so we are for now measuring just (e,e'pn).
    Will power cycle VME crate during controlled access while MCC
    is doing beam studies.

    8:30 Beam studies start.
    controlled access: Ran and BR are resetting VME crate
    and Ran is looking for N1L-26. Hall A Techs are looking into
    magnet controls

    10:00 closed the Hall, since MCC needs to send beam to Hall A
    for part of the beamstudies

    15:00 beam studies are finished, we request controlled access
    for Mark Stevens et al. to install remote reset for Q1(left)
    powersupply. (Bryan Moffit did some work in the right HRS)

    16:00 Mark Stevens/Scott Spiegel are still in the Hall
    new shift crew arrives


    Runs:
    =====

    3095, production 192k, no ADCs from ROC 5 (BigBite)
    3096, junk
    ...
    3102, still junk
    3103, cosmics (no target, but some beam in the Hall)
    3104, cosmics (no target, and no beam)
    3105, cosmics (no target, and no beam)